OpenStudy (anonymous):

Solve this system of equations by using the substitution method... y - 2x = 5 -2y + 7x = -4

jimthompson5910 (jim_thompson5910):

isolate y in y - 2x = 5 to get y - 2x = 5 y - 2x+2x = 5+2x y = 5+2x

jimthompson5910 (jim_thompson5910):

then plug this into the second equation -2y + 7x = -4 -2(5+2x) + 7x = -4 ... replace y with 5+2x (since the two are equal) now solve for x

jimthompson5910 (jim_thompson5910):

once you get x, you can find y

OpenStudy (anonymous):

K. Thanks !

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

when I plug that equation in...I got confused

jimthompson5910 (jim_thompson5910):

so you're still stuck? or you know what to do now?

OpenStudy (anonymous):

I thought I had it , then all of a sudden I got confused.. after plugging it in I got -10 - 4x + 7x = -4 ... is that right ?

jimthompson5910 (jim_thompson5910):

so far so good

OpenStudy (anonymous):

then add -10 to both sides?

jimthompson5910 (jim_thompson5910):

just 10, so add 10 to both sides

OpenStudy (anonymous):

then I get -4x+7x=6 .. right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

then I got 3x=6 and then divided both sides by 3 and x =2 , now all I do is plug in the 2 to the other equation & I get my answer ! : )

jimthompson5910 (jim_thompson5910):

yep x = 2

jimthompson5910 (jim_thompson5910):

y = 5+2x so you just plug x = 2 into that to find y

OpenStudy (anonymous):

y = 9 !

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Thank you so much !

jimthompson5910 (jim_thompson5910):

yw